### v1

# A Reinforcement Learning Approach for RIS-aided Fair Communications

Abstract
Reconfigurable Intelligent Surfaces (RISs) are composed of physical elements that can dynamically alter electromagnetic wave properties to enhance beamforming and leading to improvements in areas with low coverage properties. They have the potential to be combined with Reinforcement Learning (RL) techniques to achieve network performance and energy efficiency via optimization techniques. In addition to performance and energy improvements, it is also crucial to consider the concept of fair communications. RISs must ensure that User Equipment (UE) units receive their signals with adequate strength, without other UE being deprived of service due to insufficient power. In this paper, we address such a problem. We explore the fairness properties of previous work and propose a novel method that aims at obtaining an efficient and fair duplex RIS-RL system for multiple legitimate UE units. We report and discuss our experimental work and simulation results. We also release our code and datasets to foster further research in the topic.
Index Terms:
Reinforcement Learning, Deep Reinforcement Learning, Machine Learning, Reconfigurable Intelligent Surface, User Equipment, Security, Eavesdropping, Fairness.
I Introduction
To ensure that future Beyond-Fifth-Generation (B5G) and 6G network systems are deployable and usable, it is essential to guarantee that they meet the necessary service and security criteria for optimal operation. Reconfigurable Intelligent Surfaces (RISs), crucial and innovative elements of these futuristic networks, are subject to the same requirements due to their strategic role in the communication protocol.
RISs offer significant potential for enhancing wireless communication in 5G and B5G network systems [1]. RISs are passive sub-wavelength elements that can dynamically alter electromagnetic wave properties to perform beamforming, with the goal of optimizing network performance [2]. RISs can also enhance network coverage by redirecting signals, mitigating blockages, and extending reach in urban environments, ensuring consistent connectivity [3]. They also contribute to energy efficiency by optimizing signal paths and reducing the need for high-power transmissions, supporting sustainable operations [4]. By redirecting the signal from the base antenna to end-users through the dynamic configuration of their sub-elements, RIS operate at the physical level. Through beamforming, they are responsible for ensuring that the physical signal correctly reaches the intended user, thereby directing the signal towards the right target.
The strategic positioning of RIS, combined with the necessity for rapid adaptability to support future scalability—such as accommodating numerous and mobile users, as well as those entering and exiting the system—poses significant challenges. Additionally, the limited computational resources of RIS, which can be considered embedded components, further complicate this scenario. To address these challenges, the implementation of Deep Reinforcement Learning (DRL) is considered [5]. This approach is particularly suited for developing a controller capable of efficiently managing downlink and duplex communications for a large number of users.
Reinforcement learning Reinforcement Learning (RL) [6] is a type of machine learning where an agent learns to make decisions by performing actions in an environment to achieve maximum cumulative reward. Unlike supervised learning, RL does not rely on labeled input/output pairs but instead learns from the consequences of its actions. DRL [7] extends this concept by using deep neural networks to approximate the optimal action-value functions, enabling the handling of high-dimensional state spaces.
By leveraging DRL, the aim is to create a system that not only adapts dynamically to changing user demands but also optimizes resource allocation, ensuring robust and scalable communication solutions. Moreover, and beyond capacity and security concerns, it is also crucial to consider the concept of fairness, i.e., each user must receive their signal with adequate strength, without other users being deprived of service due to insufficient power. Several questions arise regarding the desired behavior and the methodology to achieve it.
In this paper, we address these challenges by proposing a novel framework that integrates DRL with fairness-aware mechanisms to enhance RIS-aided communication systems. We present experimental findings and simulation outcomes, and we make our code openly available to encourage further research in this area. Based on the existing literature, our approach uses DRL algorithms such as Deep Deterministic Policy Gradient (DDPG) and Twin Delayed Deep Deterministic Policy Gradient (TD3). This choice allows for a more effective comparison with the studies upon which our work is founded, enabling us to concentrate on the additional parameters that configure our decision-making agent. Thus, we refrain from utilizing advanced algorithms such as Soft-Actor Critic (SAC) [8] and Proximal Policy Optimization (PPO) [9], to ensure a consistent progression between our work and the articles we are based on. Our contribution includes a proposal of new reward functions to ensure equitable service distribution among users while maintaining high system performance and a discussion about the experimental results.
The paper is structured as follows. Section II surveys related work. Section III provides our problem statement. Section IV presents our approach. Section V presents experimental work and simulation results. Section VI concludes the paper.
TABLE I: Acronyms and Notation
Acronym Description
DRL Deep Reinforcement Learning
UE User Equipment
SINR Signal-to-Noise-plus-Interference Ratio
RIS Reconfigurable Intelligent Surface
Notation
$N$ Total number of composing elements in the RIS
$k$ Total number of User Equipment (UE) units
$D_{i}$ Downlink information rate for UE i
$U_{i}$ Uplink information rate by UE i
II Related Work
Recent work has already explored the application of DRL in the context of RIS. Liu et al. [10] survey different RIS design, including principles and opportunities to achieving high performance and dynamic functionality via optimization DRL techniques. Nguyen et al. [11] propose the use of DRL for early RIS-assisted solutions, demonstrating improved energy efficiency. These and other studies already highlight the potential of DRL in optimizing the performance and functionality of RIS in dynamic environments. The three main closest solutions related to our proposal are surveyed next.
Huang et al. [5] provide an approach to obtaining a DRL-controlled RIS capable of delivering signals for downlink communications from a Base Station (BS) to multiple UE units, without considering potential attackers. The goal is to maximize the sum of the data rates received by the users.
Peng et al. [12] provides a similar solution but expands the scope by considering duplex communications (uplink and downlink) and the possibility of passive eavesdroppers intercepting the signal and other interferences. The DRL-manipulated RIS is then expected to adapt to maximize the sum of data rates received by each UE while minimizing the data rates intercepted by eavesdroppers.
Yang et al. [13] addresses downlink communications via RIS for multiple UE units, taking into account the presence of eavesdroppers. The objective is to maximize the sum of signals received by each UE while minimizing the signals intercepted by eavesdroppers and adhering to Quality & Security of Service (QSoS) constraints. This involves ensuring that the data rates received by each UE is above a certain threshold and that the signal perceived by eavesdroppers is below a threshold.
Although the aforementioned approaches study how DRL-RIS can be trained and might incorporate a first notion of fairness as in Yang et al.  [13], a detailed study on this aspect is lacking to quantify and analyze how the system distributes power and how it can be improved to tackle this specific fairness challenge. In the sequel, we propose to focus on RIS-assisted duplex communications for mobile users, excluding the presence of eavesdroppers, to build a fairer system. By adopting the situational framework of Peng et al.  [12] for duplex communication and building upon the reflections initiated in [5], [13], and [12], we aim to explore the study of fairness more thoroughly. Our goal is to propose methods for developing a more robust system that efficiently addresses concerns of capacity and fairness. This work serves as an introduction and is followed by the inclusion of eavesdroppers and potential attackers in the considered scenarios. By first obtaining an efficient and fair duplex RIS-RL system for multiple legitimate UE units, we are able to build upon this foundation to explore more realistic cases with adversaries and additional environmental features.
III Problem Statement
Table I summarizes the main acronyms and notation used in our work. We consider a situation where a BS with $N_{t}$ transmit antennas and $N_{r}$ receiving antennas has to communicate with $k$ legitimate UE units, each one modeled with a single antenna for transmitting and receiving. The direct Line-of-Sight (LoS) communication is not possible due to an obstacle and we do not consider communications through this obstacle. The RIS, composed of $N$ elements, is deployed to allow communications between the BS and the UE for both downlink and uplink communications. The problem is supposed to be planar, for simplification.
Figure 1: Assumed RIS-DRL environment in our work.
We assume the RIS-DRL setup proposed by Peng et al. work [12] (cf. Figure 1), but slightly relax their theoretical assumptions, e.g., we assume absence of eavesdroppers. We also assume that UE units are fixed for a certain simulation episode. All other parameters and choices in Peng et al. work remain unchanged. Their component modeling, channel modeling, propagation model, interference and methods to compute the Signal-to-Noise-plus-Interference Ratios (SINRs) are used for reward computation. We focus our main change on the DRL elements. More specifically, novel reward functions and DRL algorithms (cf. Sections IV-D and V-A for further details).
Our DRL system jointly operates and picks up the $W\in\mathbb{C}^{N_{t}\times K}$ beamforming matrix at the BS and computes the diagonal phase shift matrix of the RIS $\Phi\in\mathbb{C}^{N\times N}$. We assume that the RIS introduces no power loss for the transiting signal. The main objective is to develop a system that maximizes the legitimate data rate transmitted to UE while ensuring that each user receives a minimum usable data rate.
Additionally, we seek to equitably distribute any surplus data rate among the users. The objective is to prevent any single user or group of users from monopolizing the majority of the bandwidth, thereby ensuring that others are not left with minimal resources. This optimization process is nonlinear, making it challenging to solve through traditional methods. The presence of multiple constraints further complicates the underlying problem.
IV Fairness Improvements
The objective is to develop a system that maximizes the legitimate data rate transmitted to UE, ensuring that each user receives a minimum usable data rate. Additionally, we aim to equitably distribute any surplus data rate among users. The goal is to prevent any single user or group from monopolizing the majority of the bandwidth, ensuring that others are not left with minimal resources. This optimization process is non-linear, making it difficult to solve using traditional methods. The presence of multiple constraints further complicates the underlying problem. Next, we summarize the reward functions of the two main approaches surveyed in Section II.
IV-A Common Terms
We introduce the following terms that are used in all of the studied and proposed reward functions:
$D_{i}=\log_{2}\left(1+\text{SINR}_{i}^{\text{D}}\right),$ (1)
$U_{i}=\log_{2}\left(1+\text{SINR}_{i}^{\text{U}}\right)$ (2)
Equation (1) represents the downlink data rate received by the $i^{\text{th}}$ user, and Equation (2) represents the uplink data rate emitted by the $i^{\text{th}}$ user and received by the BS. $\text{SINR}_{i}^{\text{D}}$ and $\text{SINR}_{i}^{\text{U}}$ are the signal-to-interference-plus-noise ratios at downlink and uplink, respectively, for user $i$.
After introducing these important terms, we proceed by detailing the first reward function that we consider, which is directly adapted from Peng et al. [12].
IV-B Peng et al. Reward Function
The reward function presented in [12] forms the basis of our formulation, emphasizing pure capacity across all users. For each user $i$, the secrecy rate $R_{i}$ is defined as:
$R_{i}$ $=\left[D_{i}+U_{i}\right]^{+},$ (3)
where $[x]^{+}$ corresponds to $\max(0,x)$. The Sum of Squared Rewards (SSR), final reward, is given by:
$C$ $=\sum_{i=1}^{k}R_{i}.$ (4)
We use this reward as the baseline function in our study since it focuses solely on pure capacity without considering fairness. We further refer to this reward function as the baseline reward function for the curves of the experiments.
Figure 2: General Actor-Critic architecture used in DDPG and TD3 algorithms
IV-C Yang et al. Reward Function
The reward function presented in [13] takes into account the Quality of Service (QoS) requirements associated to the users. The reward proposed in [13] is only for the downlink scenario and with the presence of eavesdroppers. We therefore adapt it to keep the terms that are relevant for our duplex case without any eavesdroppers:
$R_{QoS}$ $=\sum_{i=1}^{k}R_{i}-\sum_{i=1}^{k}\mu_{1}(p_{i}^{D}+p_{i}^{U})$ (5)
with $\mu_{1}$ a constant used to balance the utility and
$p_{i}^{D}=\begin{cases}1,&\text{if }D_{i}<\epsilon_{i}^{D},\forall i\in[\![1,k]\!],\\ 0,&\text{otherwise}\end{cases}$ (6)
$p_{i}^{U}=\begin{cases}1,&\text{if }U_{i}<\epsilon_{i}^{U},\forall i\in[\![1,k]\!],\\ 0,&\text{otherwise}\end{cases}$ (7)
where $\epsilon_{i}^{D}\in\mathbb{R}^{+}$ corresponds to the quality criteria for the downlink data rate associated to user $i$ and $\epsilon_{i}^{U}\in\mathbb{R}^{+}$ is the quality criteria for the uplink data rate associated to user $i$. We further refer to this reward function by using the QoS acronym for the experiments curves.
By introducing the thresholds $\epsilon_{i}^{D}$ and $\epsilon_{i}^{U}$, we can deal with a first notion of minimal coverage for each user. Nevertheless, this reward function may allow all users to reach a minimum service coverage if thresholds are reasonable, but once this condition is satisfied, the system is entirely free to maximize the data rate on a specific user to maximize easily the reward and thus leaving other users with minimum coverage, a situation that can have a high impact in case of real deployments.
We therefore propose a new reward system to improve the fairness among all others and still take into account the minimum service conditions.
IV-D Extended Reward Function
We propose two extensions to the reward function proposed in Peng et al.  [12] by adding, as well a QoS metric. These rewards are intended to reflect, for a designated user $i$, wether or not the system is currently ensuring the minimum capacity constraint and the resultant amount of data rate received.
IV-D1 Fairness and Quality of Service
For each user $i$ we introduce the following reward $R^{i}_{\text{Q}}$
$\quad R_{\text{Q}}^{i}=\begin{cases}D_{i}+U_{i}&\text{if }D_{i}\geq\epsilon_{k}^{D}\text{~{}and}\ U_{i}\geq\epsilon_{i}^{U}\\ U_{i}&\text{if }D_{i}<\epsilon_{i}^{D}\text{~{}and}\ U_{i}\geq\epsilon_{i}^{U}\\ D_{i}&\text{if }D_{i}\geq\epsilon_{i}^{D}\text{~{}and}\ U_{i}<\epsilon_{i}^{U}\\ 0&\text{otherwise}\end{cases}$ (8)
where $\epsilon_{i}^{D}\in\mathbb{R}^{+}$ and $\epsilon_{i}^{U}\in\mathbb{R}^{+}$ are the same thresholds for user $i$ as defined in IV-C .
We first define a new reward denoted as $R_{\text{fairness}}$
$R_{\text{fairness}}=\frac{\left(\sum_{i=1}^{k}R_{\text{Q}}^{i}\right)^{2}}{k\ \cdot\sum_{i=1}^{k}{R_{\text{Q}}^{i}}^{2}}$ (9)
which corresponds to the Jain’s fairness index [14] for the $R_{\text{Q}}^{i}$ rewards for all users $i$.
Notice that $R_{\text{fairness}}$ takes into account the fairness of the obtained rewards. It helps to better (and fairly) distribute the capacity of the system so one user is not obtaining an overly large proportion of the data rate, leading to deteriorated service for the other users.
Then, we extend the reward function to $R_{\text{fqos}}$, which feeds our DRL algorithm. $R_{\text{fqos}}$ is expressed as follows
$R_{\text{fqos}}=\left(1-\alpha\right)\sum_{i=1}^{k}R_{\text{Q}}^{i}+\alpha\cdot k\cdot\ R_{\text{fairness}}$ (10)
where $\alpha\in[0,1]$ is a hyperparameter that sets the tradeoff between pure QoS and pure fairness reward. We further refer to this reward function by using the Fairness & Quality of Service (FQoS) acronym for the experiments curves.
IV-D2 Fairness and Quality of Service Score
V Experimental Framework and Results
In this section, we present our experimental work and simulation results. We describe our experimental framework and the implemented algorithms. To ensure reproducibility and foster further research we release our code in a companion repository. [^1] https://github.com/alex-pierron/FAIR_RIS_AI The framework provides a comprehensive DRL environment designed for duplex communications between a BS and UE units, enhanced by a RIS. The code is entirely implemented in Python, leveraging the AI functionality from PyTorch. It currently supports DDPG [15] and TD3 [16] DRL algorithms.
The current version of our framework can freely be used to reproduce our results and expected to foster further research in the topics of this paper (i.e., application of DRL to RIS environments). It also provides graphical tools to evaluate how the system learns and monitors quantitative metrics to give a full overview of the system evolution during training. Some further details about the precise algorithms implemented in the framework are presented next. We also analyze in this section some sample simulation results obtained with our framework to evaluate the work proposed in this paper.
Figure 3: Downlink and uplink power patterns (see Section V-B).
V-A Implemented Algorithms
In addition to the DDPG algorithm [15] used in [12], we also implement and use TD3 [4] to better handle the instability of the environment and mitigate the residuals errors and overestimation for which DDPG is prone to. These two single-agent actor critic architectures are summarized in Figure 2. The state given as an input to the DRL system is kept unchanged from Peng et al.  [12]. The algorithms are implemented without any modifications from their respective original paper [15] [16] and only the reward functions have been modified to match our desired goals.
We evaluate and illustrate the impact of the reward systems proposed in Section IV, w.r.t. system capabilities and fairness. We present the results of conducting multiple training sequences, each utilizing different options for both algorithms and reward functions.
The environment remains consistent across all runs: the BS, positioned at coordinates $(0,0)$, communicates with the RIS located at $(20,100)$ in both directions. 2 stationary UE units are randomly distributed at the beginning of each episode within a square area defined by the $x$-axis range $[125,200]$ and the $y$-axis range $[25,100]$
Each episode consists of $250$ steps. Each training session consists of $5000$ episodes. We perform comprehensive training for each variation of the DRL algorithms (DDPG and TD3) and for each reward function detailed in IV-D.
We outline only the most critical parameters to facilitate comprehension and highlight significant changes between sequences. The learning rates are $\delta_{a}=5\cdot 10^{-3}$ for the actor and $\delta_{c}=1\cdot 10^{-3}$ for the critic. The target actor and critics employ soft updates with $\tau=1\cdot 10^{-4}$. The actor performs a gradient descent update every two steps, while the critic updates every step. We assume a replay buffer containing $200001$ samples with a batch size of $2048$. Our experiments primarily involve multiple parallel instances of the environment for roll-outs, gathering diverse experiences to populate the replay buffer. This approach allows the optimization process to sample from a broader set of experiences, minimizing correlation.
We employ $15$ parallel environments, each initialized with different internal seeds, to conduct the roll-outs. All experiments use the same master random seed to ensure reproducibility. For each experiment, we present the evolution of the average reward stored in the replay buffer during training. Since the replay buffer operates on a first-in, first-out basis, this average serves as a reliable indicator of the system’s learning progress from a reward perspective. The objective of the DRL system is to maximize the obtained reward, regardless of the chosen reward function.
To compare the different systems derived from various reward functions, we also plot the evolution of rewards obtained with each possible reward function. The reward function used for the optimization process is termed the decisive reward function, while the others are referred to as informative. By examining the three curves for each system, we can effectively compare their performances and behavior using a consistent standard. Additional metrics recorded and stored during the training process, along with the precise code used to generate these metrics.
Figure 4: Mean reward contained in the replay buffer at timestep $t$. Smoothed version with a rolling window of size $10$.
Figure 5: Mean value obtained per episode when computing the baseline reward function defined in Equation (4). Smoothed version with window size of $10$.
V-B Downlink and Uplink Power Patterns
Figure 3 illustrates the downlink and uplink power patterns generated using our simulation framework. This figure presents an example of beam patterns for both downlink and uplink scenarios in a duplex communication system with a single user equipment. The yellow curves represent the beam power patterns transmitted by the BS toward the RIS for the user in the downlink scenario. The purple curves depict the beamforming power patterns at the RIS level using the projected polar profile. It is directed toward either the user or the BS, depending on whether it is downlink or uplink communication. The green rectangle represents an obstacle blocking the initial signal from the BS, highlighting the necessity of using the RIS.
This simple example demonstrates the type of configuration we aim to achieve once the DRL controller is sufficiently optimized. For scenarios involving multiple user units, we expect to observe power patterns that precisely target the correct user units (or the BS for uplink communication) from the RIS.
V-C Simulation Results
Figure 4 illustrates the average reward in the replay buffer at time $t$. This metric helps us assess whether the system is learning and improving in terms of its decisive reward functions. Given that the replay buffer operates on a first-in-first-out basis, it provides a convenient means to analyze the learning progress of each system. However, this graph does not allow us to determine which reward system performs best, as the three possible reward functions operate on different scales. Nonetheless, it enables us to track the evolution of a specific system.
To facilitate a better comparison of algorithm performance, Figure 5 depicts the evolution of reward in Equation  (4) during training for all systems, even though it is only informative for some. We observe that while systems using this reward as their decisive reward excel at maximizing it, other systems—particularly those using the FQoS reward function defined in Equation (10) can achieve results approximately $90\%$ as effective as those attained by systems using reward in Equation (4). To easily compare the fairness of each system, we provide Figure 6 that is displaying the evolution of the Jain fairness index through the whole training process. This index is computed using Equation (9) for the decisive reward chosen per user.
V-D Discussion
A notable trend observable in our results is that systems utilizing reward function denoted in Equation (4) quickly achieves high performance from a reward perspective (around a nominal value of $6$ bps/Hz, in terms of spectral efficiency) and then experience a slight decline, yet remain within a high reward range for this particular function. Systems employing other reward functions exhibit similar behavior, suggesting that our systems are learning effectively, although most progress appears rapid due to the frequency of our measurements.
Figure 6 shows that systems employing the reward function defined in Equation (4) achieve a mean fairness of $0.6$ for both DDPG and TD3 algorithms. When the reward function defined in Equation (10) is used (as the decisive reward function), then the mean fairness rises to $0.8$. Given that the Jain fairness index here ranges from $\frac{1}{n}$ (worst case, where one user dominates) to $1$ (best case, where resources are perfectly split), with an $n$ value of $2$, the reward function defined in Equation (10) enhances fairness by nearly $40\%$, compared to the original reward function in Equation (4).
Figure 6: Mean Jain fairness index for each episode. Smoothed version with a rolling window of size $10$.
VI Conclusion
We explored the concept of fair communications with respect to DRL-enabled RIS environments. We assumed that RISs must ensure that UE units receive their signals with adequate strength, without other UE being deprived of service due to insufficient power. We addressed such a problem and explored the fairness properties of previous work. We proposed a novel method that aims at obtaining an efficient and fair duplex DRL-RIS system, for multiple legitimate UE units. We reported experimental work and simulation results. We have also released our code and datasets to foster further research in the topic. Perspectives for future work include a more exhaustive analysis of complex environments by adding mobility patterns to the users and increasing the scale of the whole system, i.e number of users or RIS elements.
Acknowledgments The work is supported by the French National Research Agency under the France 2030 label (NF-HiSec ANR-22-PEFT-0009).

### v2

@@
 This choice allows for a more effective comparison with the studies upon which our work is founded, enabling us to concentrate on the additional parameters that configure our decision-making agent.
-Thus, we refrain from utilizing advanced algorithms such as Soft-Actor Critic (SAC) [8] and Proximal Policy Optimization (PPO) [9], to ensure a consistent progression between our work and the articles we are based on.
+Thus, we refrain from utilizing state of the art algorithms such as Soft-Actor Critic (SAC) [8] and Proximal Policy Optimization (PPO) [9], to ensure a consistent progression between our work and the articles we are based on.
 Our contribution includes a proposal of new reward functions to ensure equitable service distribution among users while maintaining high system performance and a discussion about the experimental results.
@@
 $U_{i}$ Uplink information rate by UE i
+$R_{i}$ Secrecy Rate for user $i$
+$\alpha$ Balance parameter between fairness and capacity
+$C$ Sum Secrecy Rate
 II Related Work
@@
 The objective is to develop a system that maximizes the legitimate data rate transmitted to UE, ensuring that each user receives a minimum usable data rate.
-Additionally, we aim to equitably distribute any surplus data rate among users.
-The goal is to prevent any single user or group from monopolizing the majority of the bandwidth, ensuring that others are not left with minimal resources.
-This optimization process is non-linear, making it difficult to solve using traditional methods.
-The presence of multiple constraints further complicates the underlying problem.
+As stated above we want to provide a fair service among all users despite the non-linearity of our optimization problem.
 Next, we summarize the reward functions of the two main approaches surveyed in Section II.
@@
 IV-B Peng et al. Reward Function
-The reward function presented in [12] forms the basis of our formulation, emphasizing pure capacity across all users. For each user $i$, the secrecy rate $R_{i}$ is defined as:
+The reward function presented in [12] forms the basis of our formulation, emphasizing pure capacity across all users. For each user $i$, the Secrecy Rate (SR) $R_{i}$ is defined as:
 $R_{i}$ $=\left[D_{i}+U_{i}\right]^{+},$ (3)
-where $[x]^{+}$ corresponds to $\max(0,x)$. The Sum of Squared Rewards (SSR), final reward, is given by:
+where $[x]^{+}$ corresponds to $\max(0,x)$. The Sum Secrecy Rate (SSR), final reward, is given by:
 $C$ $=\sum_{i=1}^{k}R_{i}.$ (4)
@@
 The reward proposed in [13] is only for the downlink scenario and with the presence of eavesdroppers.
-We therefore adapt it to keep the terms that are relevant for our duplex case without any eavesdroppers:
-$R_{QoS}$ $=\sum_{i=1}^{k}R_{i}-\sum_{i=1}^{k}\mu_{1}(p_{i}^{D}+p_{i}^{U})$ (5)
-with $\mu_{1}$ a constant used to balance the utility and
+We therefore adapt it to keep the terms that are relevant for our duplex case without any eavesdroppers.
+We propose the following reward that we denote as QoS reward:
+$R_{QoS}$ $=\sum_{i=1}^{k}R_{i}-\sum_{i=1}^{k}\mu(p_{i}^{D}+p_{i}^{U})$ (5)
+with $\mu$ a constant used to balance the combined performances and the individual constraints for each user.
 $p_{i}^{D}=\begin{cases}1,&\text{if }D_{i}<\epsilon_{i}^{D},\forall i\in[\![1,k]\!],\\ 0,&\text{otherwise}\end{cases}$ (6)
@@
 IV-D Extended Reward Function
-We propose two extensions to the reward function proposed in Peng et al.
-[12] by adding, as well a QoS metric.
-These rewards are intended to reflect, for a designated user $i$, wether or not the system is currently ensuring the minimum capacity constraint and the resultant amount of data rate received.
-IV-D1 Fairness and Quality of Service
+We propose an extension to the reward function proposed in Peng et al.
+[12] by adding a QoS and a fairness incentived term.
+This reward is intended to reflect, for a designated user $i$, wether or not the system is currently ensuring the minimum capacity constraint and the resultant amount of data rate received.
 For each user $i$ we introduce the following reward $R^{i}_{\text{Q}}$
@@
 We first define a new reward denoted as $R_{\text{fairness}}$
-$R_{\text{fairness}}=\frac{\left(\sum_{i=1}^{k}R_{\text{Q}}^{i}\right)^{2}}{k\ \cdot\sum_{i=1}^{k}{R_{\text{Q}}^{i}}^{2}}$ (9)
-which corresponds to the Jain’s fairness index [14] for the $R_{\text{Q}}^{i}$ rewards for all users $i$.
-Notice that $R_{\text{fairness}}$ takes into account the fairness of the obtained rewards.
-It helps to better (and fairly) distribute the capacity of the system so one user is not obtaining an overly large proportion of the data rate, leading to deteriorated service for the other users.
+$R_{\text{fairness}}=\frac{\left(\sum_{i=1}^{k}R_{i}\right)^{2}}{k\ \cdot\sum_{i=1}^{k}{R_{i}}^{2}}$ (9)
+which corresponds to the Jain Fairness Index (JFI) [14] for the SR $R_{i}$ for all users $i$.
+Notice that $R_{\text{fairness}}$ takes into account the fairness regarding SR and SSR rewards because it is the effective data rate obtained in reality.
+It is supposed to help to better (and fairly) distribute the capacity of the system so one user is not obtaining an overly large proportion of the data rate, leading to deteriorated service for the other users.
 Then, we extend the reward function to $R_{\text{fqos}}$, which feeds our DRL algorithm. $R_{\text{fqos}}$ is expressed as follows
@@
 where $\alpha\in[0,1]$ is a hyperparameter that sets the tradeoff between pure QoS and pure fairness reward.
-We further refer to this reward function by using the Fairness & Quality of Service (FQoS) acronym for the experiments curves.
-IV-D2 Fairness and Quality of Service Score
+We further refer to this reward function by using the Fairness & Quality of Service (FQoS) acronym.
 V Experimental Framework and Results
@@
 [12].
-The algorithms are implemented without any modifications from their respective original paper [15] [16] and only the reward functions have been modified to match our desired goals.
-We evaluate and illustrate the impact of the reward systems proposed in Section IV, w.r.t.
-system capabilities and fairness.
+The algorithms are implemented without any modifications from their respective original paper [15, 16] and only the reward functions have been modified to match our desired goals.
+We evaluate and illustrate the impact of the reward systems proposed in Section IV in terms of system capabilities and fairness.
 We present the results of conducting multiple training sequences, each utilizing different options for both algorithms and reward functions.
 The environment remains consistent across all runs: the BS, positioned at coordinates $(0,0)$, communicates with the RIS located at $(20,100)$ in both directions.
-2 stationary UE units are randomly distributed at the beginning of each episode within a square area defined by the $x$-axis range $[125,200]$ and the $y$-axis range $[25,100]$
+2 stationary UE units are randomly distributed at the beginning of each episode within a square area defined by the $x$-axis range $[125,200]$ and the $y$-axis range $[25,100]$ and with a precision of 0.1 for each coordinate.
+Superposition is made impossible.
 Each episode consists of $250$ steps.
@@
 We perform comprehensive training for each variation of the DRL algorithms (DDPG and TD3) and for each reward function detailed in IV-D.
-We outline only the most critical parameters to facilitate comprehension and highlight significant changes between sequences.
+We outline only the most critical parameters to ease comprehension and highlight significant changes between sequences.
 The learning rates are $\delta_{a}=5\cdot 10^{-3}$ for the actor and $\delta_{c}=1\cdot 10^{-3}$ for the critic.
-The target actor and critics employ soft updates with $\tau=1\cdot 10^{-4}$.
+The target actor and critics employ soft updates with $\tau=5\cdot 10^{-4}$.
 The actor performs a gradient descent update every two steps, while the critic updates every step.
@@
 This approach allows the optimization process to sample from a broader set of experiences, minimizing correlation.
-We employ $15$ parallel environments, each initialized with different internal seeds, to conduct the roll-outs.
-All experiments use the same master random seed to ensure reproducibility.
-For each experiment, we present the evolution of the average reward stored in the replay buffer during training.
-Since the replay buffer operates on a first-in, first-out basis, this average serves as a reliable indicator of the system’s learning progress from a reward perspective.
-The objective of the DRL system is to maximize the obtained reward, regardless of the chosen reward function.
-To compare the different systems derived from various reward functions, we also plot the evolution of rewards obtained with each possible reward function.
+We utilize 15 parallel environments, each initialized with distinct internal seeds, to perform the roll-outs.
+To ensure reproducibility, all experiments share the same master random seed.
+For each experiment, we illustrate the progression of the average reward stored in the replay buffer during the training phase.
+As the replay buffer functions on a first-in, first-out basis, this average provides a dependable measure of the system’s learning advancement from a reward standpoint.
+The goal of the DRL system is to maximize the reward obtained, irrespective of the selected reward function.
+For the QoS and FQoS rewards, we set $\epsilon_{i}^{D}=0.15$ and $\epsilon_{i}^{U}=2$ for all users, with $\mu=2$. For the FQoS reward, we employ $\alpha=\frac{2}{3}$.
+To compare the different systems derived from various reward functions, we also show the evolution of rewards obtained with each reward function.
 The reward function used for the optimization process is termed the decisive reward function, while the others are referred to as informative.
-By examining the three curves for each system, we can effectively compare their performances and behavior using a consistent standard.
+By examining the curves for each system, we can effectively compare their performances and behavior using a consistent standard.
 Additional metrics recorded and stored during the training process, along with the precise code used to generate these metrics.
@@
 V-B Downlink and Uplink Power Patterns
-Figure 3 illustrates the downlink and uplink power patterns generated using our simulation framework.
+Figure 3 shows the downlink and uplink power patterns generated using our simulation framework.
 This figure presents an example of beam patterns for both downlink and uplink scenarios in a duplex communication system with a single user equipment.
@@
 V-C Simulation Results
-Figure 4 illustrates the average reward in the replay buffer at time $t$.
-This metric helps us assess whether the system is learning and improving in terms of its decisive reward functions.
-Given that the replay buffer operates on a first-in-first-out basis, it provides a convenient means to analyze the learning progress of each system.
-However, this graph does not allow us to determine which reward system performs best, as the three possible reward functions operate on different scales.
-Nonetheless, it enables us to track the evolution of a specific system.
-To facilitate a better comparison of algorithm performance, Figure 5 depicts the evolution of reward in Equation  (4) during training for all systems, even though it is only informative for some.
-We observe that while systems using this reward as their decisive reward excel at maximizing it, other systems—particularly those using the FQoS reward function defined in Equation (10) can achieve results approximately $90\%$ as effective as those attained by systems using reward in Equation (4).
-To easily compare the fairness of each system, we provide Figure 6 that is displaying the evolution of the Jain fairness index through the whole training process.
-This index is computed using Equation (9) for the decisive reward chosen per user.
+Figure 4 presents the average reward within the replay buffer at time $t$.
+This metric serves as an indicator of the system’s learning progress and its improvement concerning decisive reward functions.
+The first-in-first-out nature of the replay buffer facilitates the analysis of learning trajectories for each system.
+However, this figure does not allow a direct comparison of reward system performance due to the different scales of the three possible reward functions.
+It does, however, allow for the monitoring of individual system evolution over time.
+It is evident that all RIS-controllers utilizing the baseline and QoS rewards demonstrate learning over time, achieving significantly higher rewards compared to the initial training episodes.
+Conversely, those employing the FQoS reward exhibit limited improvement, potentially attributable to sub-optimal hyperparameter selection and reward shaping.
+To enable a comparative assessment of algorithm performance, Figure 5 illustrates the progression of the baseline reward, cf.
+Equation (4), during the training phase for all systems, despite its relevance to only certain configurations.
+From this perspective, the systems that demonstrate superior performance, without considering fairness, are those trained with the QoS, cf.
+Equation (5), and the baseline rewards.
+Notably, the system trained with the QoS reward and employing DDPG surpasses all other trained systems by at least 15%, achieving a stabilized mean SSR of approximately 8 bps/Hz by the end of training.
+In comparison, the baseline reward reaches up to 7 bps/Hz for the TD3 variant.
+Systems trained with FQoS, cf.
+Equation (10), exhibit inferior performance, achieving only 6 bps/Hz, indicating challenges in balancing performance with fairness considerations.
 V-D Discussion
-A notable trend observable in our results is that systems utilizing reward function denoted in Equation (4) quickly achieves high performance from a reward perspective (around a nominal value of $6$ bps/Hz, in terms of spectral efficiency) and then experience a slight decline, yet remain within a high reward range for this particular function.
-Systems employing other reward functions exhibit similar behavior, suggesting that our systems are learning effectively, although most progress appears rapid due to the frequency of our measurements.
-Figure 6 shows that systems employing the reward function defined in Equation (4) achieve a mean fairness of $0.6$ for both DDPG and TD3 algorithms.
-When the reward function defined in Equation (10) is used (as the decisive reward function), then the mean fairness rises to $0.8$.
-Given that the Jain fairness index here ranges from $\frac{1}{n}$ (worst case, where one user dominates) to $1$ (best case, where resources are perfectly split), with an $n$ value of $2$, the reward function defined in Equation (10) enhances fairness by nearly $40\%$, compared to the original reward function in Equation (4).
-Figure 6: Mean Jain fairness index for each episode. Smoothed version with a rolling window of size $10$.
+For an evaluation of the RIS-controllers fairness, Figure 6 displays the evolution of the JFI throughout the training process.
+This index is calculated using Equation (9) for the baseline reward, reflecting the actual signal received by users.
+The figure reveals that the RIS-controllers previously identified as achieving the highest SSR, specifically those using the QoS reward with DDPG, exhibit lower fairness indices, ranging between 0.64 and 0.61 on average.
+Given that the corresponding JFI ranges from $\frac{1}{k}$ in the most unfair scenario to 1 in a perfectly fair distribution, these values suggest that one user predominantly receives the majority of the rate.
+This occurs despite the minimum quality of service constraints being met in the case of the controller using the QoS reward with DDPG.
+Systems with lower performance achieve higher average fairness indices, between 0.67 and 0.65.
+However, this marginal increase in fairness does not compensate for the loss in of pure SSR in Figure 5.
+To further understand the fairness extremes of the trained RIS-controllers, it is necessary to examine the JFI corresponding to the maximum decisive reward obtained in each episode.
+Figure 6: Mean Jain Fairness Index (JFI) for each episode. Smoothed version with a rolling window of size $50$.
+Figure 7 shows the JFI obtained for the maximum reward value reached during the episode.
+It reveals more pronounced behavioral differences among the systems.
+Those achieving higher performances with respect to the baseline reward exhibit significantly lower fairness, below 0.6, whereas systems utilizing the FQoS maintain a consistent fairness level as previously described.
+We can conclude that, on average and at peak performance, the current systems are unfair.
+They heavily favor one user while only satisfying the minimum quality of service constraints for the other.
+This analysis underscores the importance of embedding fairness signals in reward systems.
+Although the proposed FQoS reward achieves suboptimal performance and is not without flaws, the study demonstrates that alternative rewards do not offer superior fairness.
+There is a clear need for extensive hyperparameter research with the existing reward functions, as well as a more comprehensive exploration of strategies to balance signal maximization, quality of service constraints, and fairness.
+Since we only changed the reward functions, changing the composition of the state could also leverage additional performance and represent an other research direction.
+This paves the way for enhanced RIS-controllers and ultimately lead to better and fairer B5G communication systems.
+Figure 7: Jain Fairness Index (JFI) for the best reward obtained at each episode. Smoothed version with a rolling window of size $50$.
 VI Conclusion
@@
 We assumed that RISs must ensure that UE units receive their signals with adequate strength, without other UE being deprived of service due to insufficient power.
-We addressed such a problem and explored the fairness properties of previous work.
+We investigated such a problem and explored the fairness properties of previous work.
 We proposed a novel method that aims at obtaining an efficient and fair duplex DRL-RIS system, for multiple legitimate UE units.
 We reported experimental work and simulation results.
-We have also released our code and datasets to foster further research in the topic.
-Perspectives for future work include a more exhaustive analysis of complex environments by adding mobility patterns to the users and increasing the scale of the whole system, i.e number of users or RIS elements.
+We have also released our code and experiments to foster further research on this topic.
+Future research directions may include an investigation into hyper-parameter optimization and reward engineering to achieve a better balance between pure capacity and perfect fairness, thereby attaining optimal performance.
+Subsequent studies could concentrate on the detailed analysis of more complex environments by incorporating user mobility patterns and expanding the scale of the system, specifically in terms of the number of users or RIS elements.
 Acknowledgments The work is supported by the French National Research Agency under the France 2030 label (NF-HiSec ANR-22-PEFT-0009).